OpenStudy (anonymous):

Please write y=3x-6 in the form y=Af(B(x-h))+k

OpenStudy (anonymous):

i get y=3(x-2)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

so a=3, b=1, h=2, k=0

ganeshie8 (ganeshie8):

Looks perfect!

OpenStudy (anonymous):

wait so is y=3(x-2) in the right form?

ganeshie8 (ganeshie8):

i prefer, y = 3(x-0) - 6

ganeshie8 (ganeshie8):

A = 3 B = 1 h = 0 k = -6

ganeshie8 (ganeshie8):

but your form is also right

ganeshie8 (ganeshie8):

both are right